undeadknight26 (undeadknight26):

Fiber-optic cables are used widely for wiring in aircraft and submarines. When light passes through a fiber-optic cable, its intensity decreases with the increase in the length of the cable. If 1200 lumens of light enters the cable, the intensity of light decreases by 2.5% per meter of the cable. Part A: Can this situation be represented by an exponential function? Justify your answer. (2 points) Would this mean...

undeadknight26 (undeadknight26):

F(x) = 1200(2.5)^x?

OpenStudy (math&ing001):

I think it should be F(x)=1200-1200*0.025*x with x=nbr of meters of the cable

undeadknight26 (undeadknight26):

why would it be 1200-1200?

undeadknight26 (undeadknight26):

Wait no i think im doing this wrong i think im just supposed to say yes or no if it can be a exponential function.

OpenStudy (math&ing001):

It's not 1200-1200. It decreases by 2.5% lumens per meter of the cable so it's: F(x)=1200-1200*0.025*x=1200*(1-0.025*x)

undeadknight26 (undeadknight26):

so that means it can be represented by a exponential function right?

OpenStudy (math&ing001):

Nope

undeadknight26 (undeadknight26):

How can i prove that?

OpenStudy (math&ing001):

Because it decreases periodically, i wouldn't say that it could be represented by an exponential function.

undeadknight26 (undeadknight26):

a few more questions?

undeadknight26 (undeadknight26):

How could i represent this with a function? F(x) = 1200/2.5?

undeadknight26 (undeadknight26):

i have one more after this....

OpenStudy (math&ing001):

You can use the one I pointed above

undeadknight26 (undeadknight26):

this one?F(x)=1200-1200*0.025*x wouldn't it have to follow y = mx + b though?

undeadknight26 (undeadknight26):

@math&ing001

OpenStudy (math&ing001):

It does m=-1200*0.025 and b=1200

undeadknight26 (undeadknight26):

oh ok thx.

undeadknight26 (undeadknight26):

The last question is: Some scientists are trying to make a cable for which the intensity of light would decrease by 3 lumens per unit length of the cable. Can this situation be represented by an exponential function? Justify your answer and write the appropriate function to represent this situation if 1200 lumens of light enter the cable.

OpenStudy (math&ing001):

What do you mean by "unit length" ?

undeadknight26 (undeadknight26):

im guessing the length of the cable.

undeadknight26 (undeadknight26):

Ill make a new question...

OpenStudy (anonymous):

this messed me up ... lol

undeadknight26 (undeadknight26):

*sigh* gomen.

OpenStudy (anonymous):

@undeadknight26 did you get this question correct?